if f(1)=5 and f(n)=-3f(n-1) then f(4)=
Answer by greenestamps(13327)



The statement "f(n)=-3f(n-1)" defines a sequence of numbers recursively -- that is, each term in the sequence is determined by the term(s) before it.

Specifically, in this example, it says that the n-th term in the sequence is -3 times the (n-1)th term in the sequence. In other words, each term in the sequence is found by multiplying the preceding term by -3.

Given f(1)=5 (i.e., the first term is 5), you can easily find
f(2) = -3*f(1) (second term is -3 times the first)
f(3) = -3*f(2) etc.
f(4) = -3*f(3)